# DocForge-Bench: A Comprehensive Benchmark for Document Forgery Detection and Analysis

### v1

@@ A.3 Full Results Tables
| Method | Type | DocTamper | T-SROIE | RTM | IC13 | Receipt | FantasyID | MixTamper | FSTS-1.5k | ||||||||
| AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | ||
| DocTamper | Doc | 0.995 | 0.942 | 0.811 | 0.078 | 0.660 | 0.068 | 0.748 | 0.312 | 0.731 | 0.037 | 0.754 | 0.167 | 0.940 | 0.670 | 0.877 | 0.630 |
| DTD (FH) | Doc | 0.988 | 0.786 | 0.872 | 0.106 | 0.730 | 0.092 | 0.857 | 0.335 | 0.813 | 0.046 | 0.853 | 0.218 | 0.763 | 0.204 | 0.730 | 0.347 |
| FFDN | Doc | 0.991 | 0.806 | 0.904 | 0.155 | 0.741 | 0.090 | 0.895 | 0.352 | 0.842 | 0.088 | 0.880 | 0.307 | 0.770 | 0.160 | 0.736 | 0.337 |
| CAFTB-Net | Doc | 0.995 | 0.906 | 0.810 | 0.060 | 0.704 | 0.050 | 0.845 | 0.206 | 0.791 | 0.031 | 0.796 | 0.246 | 0.965 | 0.691 | 0.885 | 0.695 |
| TIFDM | Doc | 0.994 | 0.813 | 0.806 | 0.076 | 0.719 | 0.071 | 0.788 | 0.277 | 0.762 | 0.037 | 0.879 | 0.348 | 0.900 | 0.400 | 0.792 | 0.417 |
| ASCFormer | Doc | 0.585 | 0.180 | 0.924 | 0.780 | 0.631 | 0.266 | 0.508 | 0.022 | 0.613 | 0.261 | 0.610 | 0.233 | 0.636 | 0.314 | 0.560 | 0.133 |
| ADCD-Net | Doc | 0.974 | 0.822 | 0.899 | 0.315 | 0.682 | 0.142 | 0.731 | 0.324 | 0.757 | 0.129 | 0.750 | 0.226 | 0.802 | 0.402 | 0.704 | 0.446 |
| TruFor | General | 0.762 | 0.232 | 0.978 | 0.697 | 0.663 | 0.140 | 0.782 | 0.271 | 0.743 | 0.263 | 0.830 | 0.345 | 0.988 | 0.739 | 0.916 | 0.668 |
| MVSS-Net | General | 0.585 | 0.090 | 0.613 | 0.039 | 0.550 | 0.041 | 0.668 | 0.215 | 0.592 | 0.050 | 0.647 | 0.102 | 0.804 | 0.363 | 0.546 | 0.184 |
| PSCC-Net | General | 0.770 | 0.248 | 0.629 | 0.062 | 0.573 | 0.057 | 0.665 | 0.248 | 0.708 | 0.033 | 0.735 | 0.149 | 0.913 | 0.637 | 0.832 | 0.502 |
| IML-ViT | General | 0.688 | 0.143 | 0.541 | 0.036 | 0.587 | 0.058 | 0.768 | 0.192 | 0.566 | 0.024 | 0.576 | 0.061 | 0.862 | 0.430 | 0.732 | 0.318 |
| SAFIRE | General | 0.660 | 0.163 | 0.516 | 0.029 | 0.513 | 0.045 | 0.546 | 0.191 | 0.543 | 0.023 | 0.493 | 0.065 | 0.796 | 0.499 | 0.586 | 0.243 |
| ManTraNet | General | 0.589 | 0.062 | 0.471 | 0.033 | 0.532 | 0.050 | 0.555 | 0.182 | 0.615 | 0.028 | 0.483 | 0.056 | 0.733 | 0.181 | 0.631 | 0.203 |
| CAT-Net | General | 0.982 | 0.744 | 0.860 | 0.382 | 0.624 | 0.087 | 0.625 | 0.125 | 0.781 | 0.316 | 0.835 | 0.278 | 0.990 | 0.747 | 0.880 | 0.625 |
@@ Appendix B Method Validation Against Published Benchmarks
Fine-tuning: ForensicHub [Du and others, 2025] (FFDN, DTD, TIFDM) and the original DTD paper fine-tune on DocTamper training data before evaluation; we use frozen pretrained weights. CAFTB-Net and TIFDM’s original weights were pretrained on DocTamper, so our frozen-weight numbers exceed the ForensicHub re-trained baseline (which starts from scratch).
@@
| TIFDM | DocTamper test | 0.742 | 0.259 | Du and others [2025] (NeurIPS 2025) | Retrained from scratch on DocTamper | Same explanation as CAFTB-Net: original pretrained weights vs. ForensicHub scratch retraining. |

### v2

@@ A.3 Full Results Tables
| §IML-ViT crops input to 10241024; 79% of FantasyID forgery pixels fall outside this crop, severely underestimating true localization capability. | |||||||||
| Method | Type | DocTamper | T-SROIE | RTM | IC13 | Receipt | FantasyID | FSTS-1.5k | |||||||
| AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | AUC | Opt-F1 | ||
| DocTamper | Doc | 0.995 | 0.942 | 0.811 | 0.078 | 0.660 | 0.068 | 0.748 | 0.312 | 0.731 | 0.037 | 0.754 | 0.167 | 0.877 | 0.630 |
| DTD (FH) | Doc | 0.988 | 0.786 | 0.872 | 0.106 | 0.730 | 0.092 | 0.857 | 0.335 | 0.813 | 0.046 | 0.853 | 0.218 | 0.730 | 0.347 |
| FFDN | Doc | 0.991 | 0.806 | 0.904 | 0.155 | 0.741 | 0.090 | 0.895 | 0.352 | 0.842 | 0.088 | 0.880 | 0.307 | 0.736 | 0.337 |
| CAFTB-Net | Doc | 0.995 | 0.906 | 0.810 | 0.060 | 0.704 | 0.050 | 0.845 | 0.206 | 0.791 | 0.031 | 0.796 | 0.246 | 0.885 | 0.695 |
| TIFDM | Doc | 0.994 | 0.813 | 0.806 | 0.076 | 0.719 | 0.071 | 0.788 | 0.277 | 0.762 | 0.037 | 0.879 | 0.348 | 0.792 | 0.417 |
| ASCFormer | Doc | 0.585 | 0.180 | 0.924 | 0.780 | 0.631 | 0.266 | 0.508 | 0.022 | 0.613 | 0.261 | 0.610 | 0.233 | 0.560 | 0.133 |
| ADCD-Net | Doc | 0.974 | 0.822 | 0.899 | 0.315 | 0.682 | 0.142 | 0.731 | 0.324 | 0.757 | 0.129 | 0.750 | 0.226 | 0.704 | 0.446 |
| TruFor | General | 0.762 | 0.232 | 0.978 | 0.697 | 0.663 | 0.140 | 0.782 | 0.271 | 0.743 | 0.263 | 0.830 | 0.345 | 0.916 | 0.668 |
| MVSS-Net | General | 0.585 | 0.090 | 0.613 | 0.039 | 0.550 | 0.041 | 0.668 | 0.215 | 0.592 | 0.050 | 0.647 | 0.102 | 0.546 | 0.184 |
| PSCC-Net | General | 0.770 | 0.248 | 0.629 | 0.062 | 0.573 | 0.057 | 0.665 | 0.248 | 0.708 | 0.033 | 0.735 | 0.149 | 0.832 | 0.502 |
| IML-ViT | General | 0.688 | 0.143 | 0.541 | 0.036 | 0.587 | 0.058 | 0.768 | 0.192 | 0.566 | 0.024 | 0.576 | 0.061 | 0.732 | 0.318 |
| SAFIRE | General | 0.660 | 0.163 | 0.516 | 0.029 | 0.513 | 0.045 | 0.546 | 0.191 | 0.543 | 0.023 | 0.493 | 0.065 | 0.586 | 0.243 |
| ManTraNet | General | 0.589 | 0.062 | 0.471 | 0.033 | 0.532 | 0.050 | 0.555 | 0.182 | 0.615 | 0.028 | 0.483 | 0.056 | 0.631 | 0.203 |
| CAT-Net | General | 0.982 | 0.744 | 0.860 | 0.382 | 0.624 | 0.087 | 0.625 | 0.125 | 0.781 | 0.316 | 0.835 | 0.278 | 0.880 | 0.625 |
@@ Appendix B Method Validation Against Published Benchmarks
Fine-tuning: ForensicHub [Du and others, 2025] (FFDN, DTD, TIFDM) and the original DTD paper fine-tune on DocTamper training data before evaluation; we use frozen pretrained weights. CAFTB-Net’s original weights were pretrained on DocTamper, so our frozen-weight numbers exceed the ForensicHub re-trained baseline (which starts from scratch). TIFDM’s training corpus is not publicly documented; however, its original weights achieve F1=0.742 on DocTamper versus ForensicHub’s from-scratch 0.259 baseline—a 2.9 gap consistent with in-domain pretraining, suggesting possible DocTamper overlap that we cannot confirm.
@@
| TIFDM | DocTamper test | 0.742 | 0.259 | Du and others [2025] (NeurIPS 2025) | Retrained from scratch on DocTamper | Training corpus undisclosed; the 2.9 performance gap over ForensicHub scratch retraining is consistent with possible DocTamper pretraining, but cannot be confirmed. All seven non-DocTamper datasets are unambiguously cross-domain. |